OpenStudy (anonymous):

what is the slope intercept form of the equation y-7=-5/2(x+4)?? need help!

OpenStudy (anonymous):

slope intercept form looks like y = mx + b So you need to simplify your equation until it resembles y = mx + b y - 7 = -(5/2)(x + 4) Start by multiplying out the stuff on the right side... y - 7 = -(5/2)x -(5/2)(4) y - 7 = -(5/2)x - 10 Then keep going to get the y all alone on the left.